. Grocers Inc. is considering offering a purified water service through a contract company that would locate a machine on the inside of the market store. There are two contract companies Grocers Inc. is considering, ClearWater and PureVida. ClearWater would charge an annual lease fee of $800 for set-up of the machine and for this machine, there is a utility cost of $0.10 for every gallon of water dispensed and ClearWater charges $0.05 for maintenance. For PureVida, the annual lease fee is $700, the utility cost is $0.12 for every gallon and PureVida charges $0.06 for maintenance. Grocers Inc. customers would purchase refilled gallons of water for $0.97. a. What is the annual break-even point for each option? (10 pts) b. At what volume in number of gallons would the two options have the same cost? (10 pts) c. At what forecasted volume should Grocers Inc. select ClearWater and what volume should they select PureVida and why? (5 pts)
Answer #1

BEP for clearwater = Fixed cost/(Revenue- variable cost per unit)

= 800/ (0.97-0.1-0.05)

=975.61

BEP for PureVida= Fixed cost/(Revenue- variable cost per unit)

= 700/ (0.97-0.12-0.06)

=886.08

B: Let the volume be x

Cost for Clearwater = Cost for Purevida

800+(0.1+0.05)x = 700+(0.12+0.06)x

800+0.15x= 700+0.18x

100=0.03x

X= 100/0.03 = 3333.33 units

C: If the volume is above 3333.33 units say 3334 units,

Cost of clearwater = 800+ 0.15*3334 = $1300.10

Cost of PureVida = 700+ 0.18*3334 = $1300.12

Hence the cost of Clearwater is lower beyond 3333.33 units.

Cost of Purevida is lower below 3333.33 units.